430 Carriage House Receiver Log

as of Monday, 20-Nov-2017 11:35:46 AST

Go to Most Recent Entry


NameEmailDateComment
ANS anolla@naic.edu 9/30/99 Problem:There was not cal. in the LRR # 2 system. Correction: There it was a bad line driver (8T14) in the Ch. 1 cal. Buffers port B.
ANS anolla@naic.edu 10/4/99 Problem: Receiver dewar warmed up due to power dip since .10/2/99. Correction: Receiver dewar was cooled down.
ANS anolla@naic.edu 10/5/99 Problem: There is a 10 lbs. Helium leak per week in the cryo system. Correction: It was found in the uphill position cryo hoses (the ones used to install the 21 Cm comp. in case a failure of the regular comp. of the 430 Mhz receiver)
ANS anolla@naic.edu 10/8/99 Problem: Dewar warmed up due to a power dip at 11:30 Correction: Found system warmed , cooldown started at 15:30 , ended at 17:15.
ANS anolla@naic.edu 10/13/99 Problem: No CRR receivers after ligthning and power failure. Correction: Found the CRR 1,2 & 3 amp on dewar burned out. Also the cal. driver , step att. , and monoplexer were burned out .
ANS anolla@naic.edu 10/18/99 Due to a 10 db less gain at the amplifier replaced on dewar (CRR 1 / LRR 1) the levels at the receiver downconverters were adjusted . At CRR 1 a tnc 10 db fix pad was removed , the var att. stay set at 6 db. In the CRR 2 a 10 dB tnc removed ,still with a 6 dB fixed , the var att. still set at 0 db. In CRR 3 a 6 db tnc fixed att. replaced with a 1 db att. and the var. att. was setted at 4 db instead 8 db.
ANS anolla@naic.edu 10/29/99 The cryo comp. was down due to a power dip on 10/28/99. Jrosa and Wgenter cooled it down.Also they replece the Austin comp. (ED 021) with the CTI one (ED 016).
ANS anolla@naic.edu 11/2/99 LRR 2 gain was increased by 6 db at the post amp in order to compensate it to the same level as LRR1.
ANS anolla@naic.edu 5/30/2000 There is a problem with the CRR receivers ans is that when the 430 mhz xmmter is on some kind of noise gets into the receiver causing an unstability on it.
ANS anolla@naic.edu 9/20/2000 The Cal var. att. in the CRR /LRR 1 was not working properly.JR replaced the Weinschel att on unit.
ANS anolla@naic.edu 9/28/2000 Both Dewar RF amplifiers were burned out during the T 1338 run.The first one was burned on 9/ /2000 ,the reason we don't know.The second one (LRR 2 ) was burned on 9/27/2000 , and it was caused by a faulty diode in the monoplexer.
ANS anolla@naic.edu 10/4/2000 Receiver was cooled down , LRR 1 T Sys is 74.0k A/L 6.5db & LRR 2 T Sys is 62k A/L 7.0 db.
ANS anolla@naic.edu 1/18/2001 430 Mhz Chnn. A was very unstable ,due to this condition we replaced it with chann. B amp
ANS anolla@naic.edu 1/24/2001 The CRR 1 460 Mhz L.O. became bad , it had a bad 2N3325 xsistor in the 480 mhz multiplier (second amp stage,it was replaced with a 2N2907 ). Also the final 460 mhz amp had a bad stage (third).
ANS anolla@naic.edu 1/30/2001 Removed the Chann A amp to be checked out, it was installed back in the afternoon (found 3 problems with it, broken capacitor and two cold joints)
ANS anolla@naic.edu 1/31/2001 Dewar been cooled down.
ANS anolla@naic.edu 2/1/2001 The receiver was warmed up , due that chn A still unstable. After a close inspection inside the Dewar it was found that the output rf cable of chan A had both conn with cold solder joints.
ANS anolla@naic.edu 2/2/2001 The system was cooled down, after checking the receivers, the unstability disappear. But we found during the check out that the cal step att. of LRR 2 is not switching, is stocked at the 36 K cal.
ANS anolla@naic.edu 2/5/2001 The receiver cryo was at room temp, found compressor tripped off. Restart comp and cooling down system. Also the LRR 2 Cal stepper control replaced, cal was stocked at 36 k.
ANS anolla@naic.edu 2/6/2001 LRR 2 bias Id adjusted in order to increase Gain and better N.F,was 6.58 ma and was adjusted to 5.4 ma.
ANS anolla@naic.edu 3/26/2001 CRR1/LRR1 amp burned out , JR replaced it with the LRR2 one.Also he replaced the wafer sw of the xmmter interlock ("Y" sw control) at the recvr control with a shorting kind.One of the "Y" sw is bad , VI by-passed it.
ANS anolla@naic.edu 3/27/2001 Receiver turned off in order to replace burned amp on CRR/LRR1.
ANS anolla@naic.edu 3/28/2001 CRR/LRR1 burned out amp replaced ,it was sn 4302 , the replcement is sn 4304. VI replaced the "Y" Sw motor , it had bad starting capacitor on it.
ANS anolla@naic.edu 3/30/2001 The recvr Tsys was checked .LRR 1 A/L was 6 db and the tsys. was 76 k.The LRR 2 A/L was 7.5 db and the tsys. was 53 k.
ANS anolla@naic.edu 4/2/2001 The CRR/LRR1 channel was burned again this morning. JR & ANS hooked the LRR2 amp into the CRR chann in order to be used by ionosphere group.
ANS anolla@naic.edu 4/4/2001 The Dewar was turned off in order to replace the burned amp. Notify control room.
ANS anolla@naic.edu 4/5/2001 The amp sn 4304 was fixed, it had burned out xsistor. This amp has a good recovery time, so they want it back in the system instead to replace it.
ANS anolla@naic.edu 4/6/2001 System was cooled down , LRR 1 A/L is 6 db & LRR 2 A/L is 7 db.L Wray checked it .
ANS anolla@naic.edu 4/9/2001 Dewar temp was 8.9 k & 50.9 k after cool down.
ANS anolla@naic.edu 4/17/2001 Monoplexer & Driver were replaced with the spare ones, mono was broken in one of its joints and the driver had the +15 VDC power supply bad.
ANS anolla@naic.edu 4/20/2001 HM & VI found a bend in the output section of the monoplexer; it was replaced with a new section. IT had a lousy isolation just 55 db instead 70 db.
ANS anolla@naic.edu 7/6/2001 The crosshead was purged due to contamination and there was not time to replace it. Also the CH 1 cryo monitor meter at the control room was oscillating giving false values of readings , a capacitor at the input data fixed it.
ANS anolla@naic.edu 7/16/2001 The rf amplifiers at the CH 1 Dewar were burned out during the 430 mhz dual beam run on 7/12/2001.In both the Mitsubishi transistor MFG-1412 was burned out. Chann. A sn 4304 (It stayed the same due to its good performance in the recovery time) Bias 1.5 V 6 ma Chann. B sn 4302 Bias 1.5 V 7 ma Now the spare one is the sn 4303 , it was on chnn. B
ANS anolla@naic.edu 8/27/2001 During a radar run on 8/26/2001 chann A amp was burned as well as the monoplexer. System using chann B set up.
ANS anolla@naic.edu 8/28/2001 Chann A amp replaced ( was sn 4304 , is sn 4303)
ANS anolla@naic.edu 10/17/2001 Channel A amp got burned on 10/11/2001 , it was replaced on 10/12/2001 . It was sn 4303 now is sn 4304.
ANS anolla@naic.edu 10/24/2001 The cryo system shut off in order to replace the cryopump tomorrow.
ANS anolla@naic.edu 10/25/2001 Cryopump ED 019 replaced with cryopump ED 015, system cooled down.
ANS anolla@naic.edu 12/13/2001 Amp A & monoplexer were burned yesterday , so chann B was installed instead , and monoplexer was installed this morning , it was fixed by VI last night.
ANS anolla@naic.edu 12/13/2001 Amp B & monopler were burned again , so monoplexer was repaired upstairs by EM and amplifer replaced by a Lunar amp. With Lunar amp we have an A/L sw of 3.5 db and a Tsys of a 180.
ANS anolla@naic.edu 12/17/2001 Both amps fixed , replace burned mitsubishi 1412 xstor on them.Chann A is 4304 and chann B is 4302. Chann A had no a/l fluctuation on it , maybe the Pamtech isolator ?We don't know.
ANS anolla@naic.edu 12/20/2001 After cool down the LRR # 1 has a Tsys of 72 k and a 5.5 db A/L sw, LRR #2 has a Tsys of 83 k ( too high, normaly is 60 k ) and a 5.7 db A/L sw.
ANS anolla@naic.edu 1/2/2002 System was warmed up , amps and isolators from both channels were taken down to be check due to high T sys on chann B and oscillations on it. G anesh found that isolator of chann B was bad. It was replaced with an spare one.Chann A is sn 4304 and chann B is sn 4302.Receiver was cooled down successfully.
ANS anolla@naic.edu 12/11/2002 Chann A recovery time was checked by Qihou .
ANS anolla@naic.edu 2/15/2002 JR fixed the spares amps for the system.He fixed the regular spare one (4303) plus an extra one that he founded that was tuned for 470 mhz (4301).
ANS anolla@naic.edu 5/21/2002 Cal system for both receivers ( LRR 1 & 2 ) was not working due to a faulty 8T14 line driver at the cal buffers upstairs , also it had a bad 7405 buffer on channel B.
ANS anolla@naic.edu 6/19/02 replaced chan a amp burned (sn 4304) on Thursday 6/13/02 with sn 4301 2.3v vdd , 10ma id
LSWlwray@naic.eduMonday, August 12, 2002 09:22:27Saturday Aug 10th: afternoon lightening storm and crowbar, channel A's cryogenic amplifier was burned. JR swapped the input to the dewar, radar run continued.
LSWlwray@naic.eduFriday, August 16, 2002 16:13:07Thursday Aug 15: warmed up the dewar in order to replace Channel A amplifier, which was burned during the recent radar run. Replaced amplifier with spare, and started cooldown in afternoon. Amplifiers were left off to expedite the cool-down.
ansanolla@naic.eduTuesday, August 20, 2002 11:11:20 Amp sn 4303 installed in chann A , it replaces sn 4301 that was burned during the dual beam run.
ansanolla@naic.eduTuesday, August 27, 2002 12:45:45The LRR # 2 recvr cal was not working properly , it was on all the time.The TTL line recvr for cal buffers (8T14) was high all the time.
ansanolla@naic.eduThursday, September 5, 2002 14:07:45The chann A was burned due to a power dip , xmitter was on the air causing a crowbar.JR & LWR sw to chann B but it also was burned due that the monoplexer isolation was just 33 db.Cryo system turned off this morning in order to replace both burned amps. ALSO we check the Angle Linear low level amp (room temp) and Eureka we have a 6 db A/L change and a Tsys betewen 95 to 100 k.
aansanolla@naaic.eduFriday, September 6, 2002 10:54:57MONOPLEXER DID IT AGAIN!!!!!!!BURNED!!!!!!!!
ansanolla@naic.eduFriday, September 6, 2002 15:16:52Dome monoplexer at CH # 1 , using the old mono driver.Plus Angle Linear amp at room temp in the CRR system.
ansanolla@naic.eduFriday, September 13, 2002 08:52:28On 9/11/2002 at AST 14:00 there was a power dip due to lightning . Due to these power dip the calibration buffers for the CRR/LRR system became inoperative (TTL line driver 8T14 short out).Also there was no A/L change in the receivers due that the post amp power supply was off ,why ???????
ansanolla@naic.eduMonday, September 16, 2002 13:29:48When checking the rf amps wiring inside the Dewar for some reason ocurre a short at the heaters wiring , it causes the heater wiring to open.
ansanolla@naic.eduThursday, September 26, 2002 14:45:11The CH 1 monoplexer was installed back after a lab test on it , but it did not work properly . So the Dome monoplexer was installed on CH 1.
ansanolla@naic.eduThursday, October 3, 2002 11:18:15The LRR 1 calibration Directional Coupler was brocken (unsolder), system having problems with its cal.
ansanolla@naic.eduThursday, October 10, 2002 13:59:58The monoplexer got burned out. RF amp ok.
ansanolla@naic.eduMonday, October 28, 2002 09:47:42New Berkshire amps on it!!!!!!! Also the cryocomp was replaced due to a failure on the existing one (now has the Austin 125 cryo comp ,good luck!!!!!!!
LSWlwray@naic.eduThursday, November 7, 2002 17:16:39while installing monoplexer and driver for a transmitting test, noticed that the amplifier in the dewar is reading overly-high current (it may be damaged). the dewar was warmed up, and will be operated on. for now, the room temperature amplifiers have been installed.
ansanolla@naic edu.Monday, November 11, 2002 12:44:48System running with an Angle Linear amp for the dual beam run. Dewar amps are cool but bias PS is off.
ansanolla@naic edu.Wednesday, November 13, 2002 14:02:15Amp was replaced , there was not A/L change on it . But the one replaced is good , so what's going on ? Second time that it happened.
ansanolla@naic edu.Wednesday, November 13, 2002 15:57:33The amp that was replaced was # 3. It checked ok at the lab.What it is , we don't know.
ansanolla@naic.eduFriday, November 15, 2002 13:18:04CH # 1 430 mhz system back to normal , it was for dual beam run using the Angle linear amp. Tsys for the An/L amp was 115k. Dewar hooked back , Tsys for LRR 1 was 55k, LRR 2 , 65k.
LSWlwray@naic.eduFriday, November 22, 2002 17:49:24Nov 21: monoplexer brought down to lab for testing. bypass put in its place.
ansanolla@naic edu.Tuesday, November 26, 2002 11:29:39System set for dual beam run. T sys was 142k. Angle linear amp # 3. New amp holding bracket on it.
ansanolla@naic edu.Wednesday, November 27, 2002 13:21:36System back to RA, cooled receiver, monoplexer is still in . Tsys was 80k for chann A & 75k for B.
ansanolla@naic edu.Monday, December 2, 2002 13:58:36System for dual beam run.Almost impossible to work on CH , it was at 20.
ansanolla@naic.eduThursday, December 5, 2002 15:29:40System back to RA .
ansanolla@naic edu.Thursday, January 2, 2003 13:02:46System set for radar.Room temp Angle Linear amp on it.
ansanolla@naic.eduFriday, January 3, 2003 15:00:15System back to Dewar (cooled amps , Berkshire ones), also the monoplexer was replaced with the one of the Greg Dome.
ans anolla@naic edu.Tuesday, January 7, 2003 12:58:56Biases for the Berkshire amps at the Dewar are: Chann A Vd Id 1 2.4 v 6.7 ma 2 2.35 v 8.1 ma Chann B Vd Id 1 1.08 v 4.1 ma 2 2.47 v 8.1 ma
aansanolla@naic edu.Thursday, January 9, 2003 11:47:06Coax cable 15 section between lower rotary joint and CH 1 was replaced with an spare one , it had water on it. Been used for 430 mhz xmmter Prepulse J1.
ans anolla@naic edu.Monday, February 3, 2003 13:26:35System was set to Radar mode since 1/28/2003 , returned to normal mode (cool down Berkshire amps , was on Angle linear amps at room temp )
ansanolla@naic edu.Wednesday, February 26, 2003 11:24:01Cryocomp was shut off by maint people working in the CH on 2/24/2003. JR/IC pumped the Dewar and restarted the cryocomp.
ansanolla@naic.eduMonday, March 3, 2003 14:23:56The 8 diode monoplexer ( Dome monoplexer) installed at the recevr in order to increase isolation.
ansanolla@naic edu.Monday, March 10, 2003 11:49:23Heavy rfi on all receivers.CRR/LRR.
ansanolla@naic edu.Monday, March 10, 2003 13:14:29RFI coming from tv station channel 14.
lswlwray@naic.eduThursday, March 13, 2003 14:28:59RFI from TV station leaking through channel B due to LNA being close to oscillation. Changed bias values FROM: Stage 1, 2: 1.053V/4.05mA, 2.147V/8.09mA TO: 1.053V/5.09mA, 2.147V/5.64mA. No more TV station leaking through.
lswlwray@naic.eduThursday, March 13, 2003 17:06:43Further tests show that the TV station is leaking through still. Now more in A than B.
ansanolla@naic.eduMonday, March 17, 2003 14:30:19Ganesh adjsuted the bias for channel B Vd1= .991 v Id1=4.05 ma Vd2= 1.964 Id2=5.50 ma
ansanolla@naic.eduWednesday, March 19, 2003 13:48:17Ganesh found that the 427 mhz rfi signal at the CH 1 430 mhz receiver is due to the intermodulation of tv station chan 14 & tv station chan 22.It is suspected that it happened since chann 14 increased its power causing these intermode problem.
ansanolla@naic.eduWednesday, March 19, 2003 14:49:08Due to a series of power dips the cryocomp shut off.Was caugth on time.
lswlwray@naic.eduWednesday, March 26, 2003 16:23:54Added a notch filter before the dewar, Ch B, to cut out 525 MHz. Ch B's Tsys is ~80K now, but the birdie at 427 is now gone. Ch A still has the 427 birdie.
ansanolla@naic.eduThursday, March 27, 2003 11:44:32The 430 Mhz channel B biases were readjusted: was: 1 stage 2nd stage Vd .99 v 1.97 v Id 4.1 ma 5.6 ma Is : Vd 1.07 v 2.47 v Id 5.03 ma 8.1 ma
ansanolla@naic.eduThursday, March 27, 2003 11:47:53Forgot to mention chnn. A biases. 1st tage is Vd 2.41 v , Id 1 is 6.67 ma. 2nd stage is Vd 2.47 v , Id is 8.1 ma
lswlwray@naic.eduThursday, March 27, 2003 16:19:51Testing the system this morning showed the TV birdies back in both channels. As the telescope tracks, they change power. Also found more birdies at 426.25, BW <1 kHz (both chanels). A comb centered at 426.832 MHz with 16 kHz spacing, each comb is <1 kHz in BW.
ansanolla@naic.eduFriday, March 28, 2003 11:52:13Bias for channel A were changed;were stage 1 (Vd 2.41 v ) (Id 6.68 ma),stage 2 (Vd 2.39 v)(Id 8.09) They are now stage 1(Vd 3.0 v) (Id 8.2 ma), stage 2 (Vd 3.0 v) (Id 10.0 ma)
ansanolla@naic.eduFriday, March 28, 2003 13:47:34Channel A biases adjusted again; is stage 1 Vd 2.43 , Id 7.1 ma. Stage 2 is Vd 2.46 v ,Id 8.6 ma.
ansanolla@naic.eduWednesday, April 9, 2003 13:57:01Add cavity filters to LRR 1 & 2 in order to reject 427 mhz from intermod from tv channel 14.
ansanolla@naic.eduWednesday, April 9, 2003 15:01:11The LRR 1 & 2 Tsys were measured after adding cavity filters. The LRR 1 gave a Tsys of 72 , it was 2.2 db with 47.5k cal and 3.7 db with 98.9k cal. The LRR 2 gave a Tsys of 65k , it was 1.9k with 36k cal and 3.1 db with 65k.
ansanolla@naic.eduFriday, April 11, 2003 09:52:42Due to a crownbar the LRR 1 receiver got burned out yesterday in the afternoon , so amp was replaced with an Anglelinear amp at room temp . It's set up is monoplexer ,Anglelinear amp , cavity filter , post amp.
ansanolla@naic,eduFriday, April 11, 2003 16:17:14CRR/LRR 1 has a new room temp Anglelinear amp. It's Tsys is +/- 80k.It's A/L change is 6 db.
lswlwray@naic.eduMonday, April 14, 2003 16:36:11warmed up dewar, plan to change burned out amplifier (ch A) tomorrow. Also, replaced commercial post amp with Angle Linear high compression amps in both channels.
ansanolla@naic.eduThursday, April 17, 2003 11:23:13The 430 mhz system has now the Angle Linear amps in both channels . Dewar is off for a while (Bershire amp on chann A was burned due to a crowbar on 4/11/2003) . LRR 1 A/L is 6db and it's Tsys is +/- 70k . Meanwhile LRR 2 A/L is 7 db and it's Tsys is +/- 75k.
ansanollla@naic.eduThursday, May 1, 2003 09:35:43Ganesh requested a receivers perfomance after the Xmmiter run ; the numbers are the following ones. System Cal Values A/L Cal T Sys LRR1 47.5/98.9k 5.7 db 2.1/3.6db 76/77k LRR2 36/65k 7.0 db 1.7/2.8db 75/72k
ansanolla@naic.eduMonday, June 23, 2003 13:02:03Biases were readjusted , they are: Chann Vd1 Id1 Vd2 Id2 A 2.51v 7.51 ma 2.00 v 6.2 ma B 1.5v 3.6 ma 1.5v 6.3 ma
ansanolla@naic.eduTuesday, June 24, 2003 13:07:06System running a dual beam run.
ansanolla@naic.eduTuesday, July 1, 2003 10:17:34System set for dual beam run until next Monday.
ansanolla@naic.eduWednesday, July 23, 2003 09:16:00System set for dual beam run. Tsys with Angle Linear amp was 105k. At midnight when swapping experiments the Angle Linear amp got burned . Ed Castro came to replace it , he used the LRR 2 amp as replacement.
ansanolla@naic.eduMonday, July 28, 2003 09:04:18On Saturday 26 the Angle Linear amp 2 got burn by a crowbar , it was replaced by amp 5 .But after its installation the Tsys stared jumping from 150k to 800k. Why ??????? We don't know.
ansanolla@naic.eduMonday, July 28, 2003 15:52:19System back to normal. Chann A Tsys is 72k and B 80k. Angle Linear amp 1 on A and 4 on B.
ansanolla@naic.eduFriday, August 1, 2003 14:09:26There is an spare Angle Linear amp (amp 3 , low) at the plastic box next to the monoplexer. The exellent limiter is with it.
ansanolla@naic.eduMonday, August 4, 2003 14:28:44The Angle Linear amp #1 (model 440GST /.33db / sn 12745 )got burned on 8/2/2003.Was replaced with amp #3.
ansanolla@naic.eduFriday, August 8, 2003 15:43:34The new 430 mhz Dewar cryo got highly contaminated when was cooled at the cryo shop.The cryopump (ED 031) was cleaned ,it had a lot of contamination on its displacers. We were using the new CTI cryocomp to be used at the CH # 1. The system is going to be run thru the absolute filter.
ansanolla@naic.eduMonday, September 22, 2003 11:17:38System set for dual beam run.
ansanolla@naic.eduTuesday, September 30, 2003 15:30:32The Angle Linear amplifier for the CRR1/LRR1 was replaced with a low level , it had a high level . CRR/LRR1 T sys is 72k and LRR 2 is 73k. It was running a dual beam.
ansanolla@naic.eduTuesday, October 7, 2003 11:37:02The CRR/LRR 1 cal driver was burned out by ligthning . The 8T14 at the cal buffer burned.
ansanolla@naic.eduTuesday, October 21, 2003 13:20:31System set for dual beam run. Spare amp at a plastic box on waveguide.
ansanolla@naic.eduMonday, October 27, 2003 13:18:13The 6 diode monoplexer from CH # 1 was replaced with the 8 diode one from the Dome. Also High level amp replaced with a low level one.
lswlwray@naic.eduThursday, October 30, 2003 16:08:16vacuum pulled on new dewar, and cooldown started at 11am.
ansanolla@naic.eduFriday, October 31, 2003 12:42:06The new Dewar was cooled it has 14.4 at 16stage and 31,7 at the 70 stage. VPG was 4.
lswlwray@naic.eduTuesday, November 4, 2003 13:44:23Amplifier 450-11 is in Ch A. 16K Bias Settings: 1.51V/4.1mA, 1.5V/5.9mA. 450-10 is in Ch B. 16K Bias Settings: 1.51V/4.3mA, 1.51V/6.1mA
ansanolla@naic.eduThursday, November 6, 2003 11:54:23# 450-10: currently in new Carriage house Dewar, Ch B Repair dates: Aug 1996, Aug 2001, Aug 2002 16K Bias settings: 1.51V/4.3mA, 1.51V/6.1mA 16K Performance: 24.8 dB gain, 1.9K at 425 MHz 295K Bias settings: 2.01V/10.3mA, 2.02V/10.4mA 295K Performance: 19.8 dB gain, 38.1K at 425 MHz # 450-11: currently in new Carriage house Dewar, Ch A Repair dates: Aug 1996, Aug 2001, Aug 2002 16K Bias settings: 1.51V/4.1mA, 1.5V/5.9mA 16K Performance: 23.7 dB gain, 1.7K at 425 MHz 295K Bias settings: 2.01V/10.4mA, 2.0V/10.2mA 295K Performance: 18.6 dB gain, 48.0K at 425 MHz
ansanolla@naic.eduMonday, November 10, 2003 14:02:35New rf set up , now the cavity filters and new post amps are inside the rf cabinet.New setting has 6 db more gain than before.
ansanolla@naic.eduTuesday, November 11, 2003 13:00:36Dual beam run , amp # 1 (low level) is on it. Amp #5 (hi level one) is a spare one , it is in the shoe box .Also there is a low level (amp # 2 ) as second back up.
ansanolla@naic.eduMonday, November 17, 2003 15:51:20Dual beam finished yesterday.
ansanolla@naic.eduWednesday, December 3, 2003 16:26:29At last the new Dewar was installed.Its A/L deflection for LRR #1 is 6.9 db and it's tsys is 61k. For LRR #2 A/Ldeflection is 7.5 db and it's tsys is also 61k. We swap the post amp in order to have attenuation control in the LRR #1.
ansanolla@naic.eduThursday, December 4, 2003 09:20:08WOW!!!!!!! Taking the readings for the new NRAO amps in CH #1 gave us the following results : LRR #1
ansanolla@naic.eduThursday, December 4, 2003 09:26:56WOW!!!!!!! Taking the data for the new NRAO amps on CH #1 gave us the following values: A/L Cal Tsys LRR #1 7.9 db 3.0 db(47.5k) 47.2K LRR #2 8.7 db 2.5 db ( 36k) 45.6k
ansanolla@naic.eduMonday, December 8, 2003 13:37:06Channel A was set with the Angle linear amp on 12/5/2003 due that they plan to use the 430 xmitter during the weekend. So these morning we set it again for RA .The A/L for chann A was 7 db Tsys 55k and for chann B was A/L 7.5 db and Tsys was 60k. LST time was 18:05.
ansanolla@naic.eduWednesday, December 10, 2003 13:23:56Channel A amp LED still flashing , it is suspected that there is a faulty ground cable iinside the Dewar.
ansanolla@naic.eduTuesday, December 16, 2003 14:12:01System set for dual beam . Angle Linear amp #1 on it, Tsys is 100k. Tsys with NRAO was 75k. There are two spares at the Carriage house,amp #2 (low level) and amp #5 (high level).
ansanolla@naic.eduTuesday, December 30, 2003 13:12:49System back to R.A. Also JRV repair the bias chassis , it had an oscillation on it making the amps LED to flicker.
ansanolla@naic.eduWednesday, December 31, 2003 10:08:45Ganesh found that the intermod signals at 430 mhz passband are 20 db higher on chann A than B, so they are seen on these channel and not on B. Also he found that hour internal heater for the RF cabinet temp control was causing a broadband noise around every minute , it was turned off.
ansanolla@naic.eduWednesday, January 14, 2004 12:53:53System was swap to radar yesterday.Today when we were swaping to RA found the cal coupler broken. WR is building a new one.
ansanolla@naic.eduThursday, January 15, 2004 16:28:47There is a new cal directional coupler in the system . It is -34 db coupling.Have to be calibrated.
lwraylwray@naic.eduThursday, January 15, 2004 16:30:37Crack found Jan 13th in the directional coupler used to input the cal signal in Ch A. Crack probably due to slight wiggling when monoplexer is removed and re-installed. Coupler brought down and fixed by Wilfredo. Tested and compared with Oct 2002 and Septemebr 1994 data to show that the S parameters have not changed by much (.10 dB max). Reinstalled on Jan 14th in the morning. Receiver shows good tsys ~80K
ansanolla@naic.eduTuesday, January 20, 2004 15:57:14Turnstile probes cleaned and adjusted to improve isolation and impedance on system.Big probe still at 9 1/4" as it was , center one was adjusted to 7 5/8" (it was 5 3/4") , also large short was push 1/2".
ansanolla@naic.eduTuesday, February 3, 2004 09:02:22Checking the LRR system on 2/3/2004 at LST 16:58 gave me the following results. A/L Cal Defl/cal Tsys LRR 1 6.7 db 2.4(47.5) 64K 4.0(98.9) 65K LRR 2 7.5 db 2.0(36.0) 61k 3.2(65.0) 59k
ansanolla@naic.eduMonday, March 15, 2004 16:14:54The rf gating sw was installed , it has +/- 30 db lost.
ansanolla@naic.eduWednesday, March 24, 2004 12:57:50 Amp A ( CRR1/LRR1 )burned out last nigth due to power glitch at midnigth. Need replacement. Dewar turn off. ANS
ansanolla@naic.eduThursday, March 25, 2004 16:01:15Channel A amp replaced with sn 450-13 it was 450-11. It's bias settings are : Vd1 1.5 vdc , Id1 3.9 ma. , Vd2 1.5 vdc, Id2 5,4 ma. It was burned at the 430 mhz on 3/23/2004.
ansanolla@naic.Monday, April 5, 2004 11:20:10Dual beam run since 3/29/2004.Cooled Tsys was chnn A 64k & B was 70k. At room temp was with amp #1 was 109k. Today was back to normal and it's T sys is , chann A 65k and B was 58k.
ansanolla@naic.eduMonday, April 19, 2004 13:13:39Channel A got burned on 4/16/2004 during a 430 mhz trsnmmiter run. It was a cooled amp and it was replaced by a room temperature one.
ansanolla@naic.eduMonday, April 19, 2004 16:24:00Channel A rf amp (450-13, cooled one) was burned out during a radar run on 4/17/2004 AST 18:30. It was replaced by a room temp one (sn 12746).
ansanolla@naic.eduWednesday, May 12, 2004 16:13:51450-11 cooled amp returned from NRAO.Was burned on 3/25/2004 run.
ansanolla@naic.eduMonday, May 17, 2004 12:50:42Dual beam run since 5/17/2004 until 5/21/2004. Amp is # 2.
ansanolla@naic.eduThursday, May 27, 2004 10:27:37Channel B swap to room temp amp in order to lower down the Dewar to replace the channel A cooled amp that was burned on 4/17/2004.Also a limiter will be added to channel A in order to protec it from 430 xmmiter crowbars.T sys are chan A 70 K deg and chan B is 80 K deg at room temp.
ansanolla@naic.eduFriday, May 28, 2004 11:53:41Dewar at cryo shop in order to replace burned rf amp on chann A . Also to try to install a limiter at the input of these chan to protect the rf amp from 430 xmitter burst.Room temp amps up for a while.
lswlwray@naic.eduTuesday, June 8, 2004 11:54:02Replaced burned amplifier 450-13 with 450-11. Current allocations: Ch A: 450-11 Ch B: 450-10
lswlwray@naic.eduThursday, June 10, 2004 16:20:40Brought receiver up to the receiver testing shack. Started the cooldown.
lswlwray@naic.eduTuesday, June 15, 2004 08:19:57Dual beam installation day... doesn't affect 430 CH, since it is already on room temperature amplifier. The dewar is down at the receiver test shack. It is cold and we're doing recovery time tests on it.
lswlwray@naic.eduMonday, June 21, 2004 09:00:18June 18: Upstairs chassis R119, postamp that feeds CRR2 and CRR3 failed. Miteq AU-2A-0550. Replaced with Angle Linear HSE-34. Tested Angle Linear amp for 1 dB compression: Pout +15 dBm, Miteq compressed earlier at +8 dBm. Similar gains (+20 dB). Miteq has broad frequency response (5-500 MHz min), but Angle Linear has narrow response, peaked around 340 MHz at +23, then a sharp slope from there.
lswlwray@naic.eduMonday, June 28, 2004 13:57:27June 22: R107 upstairs, replaced Optimax 20 dB IF (30 MHz) amplifier in CRR1/LRR1 with a Q-bit 500 (20 dB).
ansanolla@naic.eduSaturday, July 3, 2004 23:40:35I.F. compensating amp for CRR/LRR 2 at the radar room had a bad 10 db SMA att bad. It was dropping its signal for 30 db.
ansanolla@naic.eduThursday, July 8, 2004 12:47:11Radar receiver calibrator was not working properly ,both channels were bad (LRR 1 & LRR 2). The downstairs control had the line driver bad (IC 12 / MC 3487 ). Also upstairs the line receiver ( IC 4 , 8T14) was bad.
lswlwray@naic.eduFriday, July 16, 2004 11:40:37In Lab, replaced Ch A 450-11 with the suspected bad one 450-13 for Ch A. Ch B is still with 450-10. 450-13 performed well both room temp and cold. Teff and Gain are fine, and the compression testing showed no difference between Ch B and A. Left 450-13 in Ch A and the receiver was brought up to the CH on Monday July 12 although not installed... as we are waiting for maintenance to move our compressor back, they did some work to the ceiling of the carriage house.
ansanolla@naic.eduWednesday, July 21, 2004 08:26:42Dewar installed upstairsn since 7/11/2004 , been pumped since.We were not able to cool it down because maintenance people told us to remove the cryocompressor from it's place in order to give preventive maintenance to the floor in the cryocomp area.
ansanolla@naic.eduTuesday, August 17, 2004 09:22:55JRV found that the CRR/LRR 1 Jennings sw was malfunctioning , it was not making good rf contact.So the LRR 2 used as spare.
ansanolla@naic.eduThursday, August 19, 2004 13:35:05JRV bypassed gating sw due to an oscillation on the CRR 1 system (was main suspect).
ansanolla@naic.eduFriday, August 20, 2004 15:37:13There is a Dow Key relay replacing the Jennings sw in the LRR 2 receiver , Tsys is expected to be 10 deg k worst than before. The jennings was taken from these recvr in order to replace the one in LRR 1 that became bad.Also the RF gating sw was by-pass in LRR 1.
ansanolla@naic.eduTuesday, September 7, 2004 15:03:15CRR 3 receiver been used for John Mattews experiment.An RF gating sw added to it's input.
ansanolla@naic.eduFriday, September 10, 2004 11:19:30System set for dual beam run from 9/10/2004 to 9/162004.The CRR 3 was set for John Mattew experiment. Also this morning when doublechecking CH 1 receivers they were dead. It was due to a blown fuse at the L.O. driver, chassis 3-4 at rack 3 RR.Also the HP rms meter #1 is oscillating.
ansanolla@naic.eduFriday, September 10, 2004 14:36:44The CRR 3 receiver for John Mattew experiment had a ripple on the pass band.It was prodeced by its own L.O.Replace L.O. to another port (400 mhz).
ansanolla@naic.eduThursday, September 23, 2004 13:20:00Jennings A/L relay for the CRR'S/LRR 1 full of water. Was replaced with LRR 2 for a while.
ansanolla@naic.eduMonday, September 27, 2004 11:34:19Dewar been cooled.System still in the room temp Angle Linear amps.
ansanolla@naic.eduMonday, September 27, 2004 14:47:20The LRR #2 Jennings A/L sw that was full of water was cleaned.It's isolation above 60 db , insertion loss is .05 db or better.Ready to be installed back in the LRR #2 system.
ansanolla@naic.eduThursday, September 30, 2004 14:11:31System back to cooled amps. Tsys for channel A (LRR1) is 76 k deg and A/L is 6.0 db. For channel 2 (LRR 2) is 61 deg and A/L is 7.0 db. In A channel our friends are back (intermods).
ansanolla@naic.eduMonday, October 11, 2004 12:35:12System warmed up due to power failure on 10/9/2004.Have to fix loose rf connectors on Dewar.
ansanolla@naic.eduMonday, October 11, 2004 15:32:18System been cooled. RF connectors were tight.
ansanolla@naic.eduTuesday, October 12, 2004 13:22:45System at channel A on room temp amp , at channel B cooled amp on Dewar.JRV & BG found loose rf amps connectors on it.Were tight up.
ansanolla@naic.eduThursday, October 21, 2004 14:15:18CRR 3 receiver back to normal , it was set for John Mattews with external cross dipole.
ansanolla@naic.eduFriday, October 22, 2004 10:09:45When returning CRR 3 back to normal I forgot (ANS) to return the rf gating sw pulse back.It was saturated , Ganesh noticed it yesterday.
ansanolla@naic.eduFriday, November 12, 2004 12:49:27Channel A was hooked up using a limiter to the cooled amp.It's Tsys is 105 k . At room temp amp was 150 k.
ansanolla@naic.eduWednesday, November 17, 2004 09:44:10The A/L & Tsys for the LRR recvrs are: LRR 1 A/L 5.5 db Tsys 80 k LRR 2 A/L 7.0 db Tsys 60 k Note: Lot of rfi in the LRR 1 system.
ansanolla@naic.eduWednesday, November 24, 2004 09:52:47The LRR 1 A/L= 5.5 db Tsys 76k , LRR 2 A/L 6.5 db Tsys 75k.
ansanolla@naic.eduFriday, December 3, 2004 09:13:34The CH 1 readings are : CRR/LRR 1: A/L= 5.9 db Tsys=72k LRR 2: A/L= 7.0 db Tsys=57k
ansanolla@naic.eduMonday, December 6, 2004 15:29:19Dual beam been running with Dewar.Limiter between monoplexer & Dewar.
ansanolla@naic.eduFriday, December 10, 2004 15:56:19CRR 3 receiver is set for John Mathews run.
ansanolla@naic.eduThursday, December 16, 2004 09:20:33CH #1 receivers readings are: CRR/LRR1 A/L 5 db Tsys 80 k . LRR 2 A/L 7 db, Tsys 65k.
ansanolla@naic.eduTuesday, December 21, 2004 10:04:04The CRR's/LRR 1 was very unstable due to the TV stations intermodes ; in order to check it was shifted in frecuency by 5 mhz (435 mhz with LRR 1 system). Readings were as follow: CRR/LRR 1 A/L 5 db , Tsys 90k. For the LRR 2 A/L was 7 db , Tsys 70k.
ansanolla@naic.eduThursday, December 30, 2004 09:22:32430 Mhz receivers performance was at LST 15:00 the following: CRR'S/LRR1 , A/L 5.6db and Tsys was 76k. For the LRR 2 was A/L 6.9db and Tsys 65k.There were a strong intermode signals.Today we are going to try to install the cavity/notch filters to avoid the intermodes.
ansanolla@naic.eduThursday, December 30, 2004 13:47:19Cavity/Notch filter installed in the CRR'S / LRR 1 system killing the intermodes.
ansanolla@naic.eduTuesday, January 11, 2005 14:43:42The Tsys for the 430 Mhz CH 1 receivers on 1/4/2005 were as follow: CRR'S/LRR1 A/L 5.7 db and Tsys 76/79 k. For the LRR 2 were A/L 6.9 db and Tsys 65/65 k.
ansanolla@naic.eduWednesday, January 12, 2005 13:59:28The 430 Mhz CH 1 recvrs are the following for this week: CRR'S/LRR1 A/L (ant/load) is 6.0 db / Tsys is 76k.The cavity / notch filter is clening up the bandpass.Sometimes you are able to see an intermode , but very rearly.
ansanolla@naic.eduWednesday, January 12, 2005 14:02:27forgot to mention the A/L & Tsys for LRR2 this week. They are A/L 7.0 db / Tsys is 59k.There were some intermodes at the bandpass , system still witout the cavity/notch filter.
ansanolla@naic.eduThursday, January 13, 2005 12:57:58Cryocomp was charged to 270 , it was 240. Its cryotemp were 11.3 & 39.0. Cab was ok.He tank was at 2.0 k.
ansanolla@naic.eduWednesday, January 19, 2005 13:33:18The 430 Mhz measurments for this week wer: CRR'S/LRR1 A/L= 6db / Tsys 76 k deg. LRR 2 A/L 6.5 db / Tsys 60 k deg.
ansanolla@naic.eduTuesday, January 25, 2005 12:58:57The 430Mhz system measurments for these week were the following ones: CRR's/LRR1 A/L 6.0 db , Tsys 76k. The LRR 2 was A/L 5.5 db , Tsys 90k.
ansanolla@naic.eduWednesday, February 2, 2005 13:31:56Cavity filter installed in the LRR 2 side. For some reason system A/L swap is 9 db instead 7db and Tsys improved from 60/65 k to 40/45 k. These happened when the Jennings sw was replaced with a Dow-Key relay around a week ago.
ansanolla@naic.eduThursday, February 3, 2005 11:07:19There it was a problem with the LRR 2 measurments , too good to be true.It was that the A/L relay was bad , no contac at the ant side. So was replaced and everything back to normal. The relay is the Dow-Key one.The readings for these week are: CRR'S /LRR 1 A/L 5.4 db ,Tsys 80k. LRR 2 A/L 6.0 db , Tsys 85 k.
ansanolla@naic.eduTuesday, February 8, 2005 15:47:28The IF cables for the 430 mhz receivers were replaced between the upper rotary joint & lower rotary joint.The CRR/LRR 1 if cable had the shield broken.Also there is a spare cable in case of emergency. Cables are 52' long (RG-214 was used instead RG-9).
ansanolla@naic.eduTuesday, February 8, 2005 15:54:54Forgot to mention that the CRR 2 L.O. is bad .Replaced with a 400 Mhz L.O. for a while.LSW working on L.O. unit.
ansanolla@naic.eduThursday, February 10, 2005 16:30:52Crr 2 L.O. hooked up back. Had a short and bad power supply.
ansanolla@naic.eduWednesday, February 16, 2005 14:56:10The cal driver buffers were modified to be driven with fiber optic link.
ansanolla@naic.eduFriday, February 25, 2005 16:25:07The Jennings A/L sw on the CRR/LRR1 system replaced with the Dome channel B one . It is intermmitant.
ansanolla@naic.eduMonday, February 28, 2005 14:34:02The LRR 2 post amp #2 and BP filter (430/10 Mhz BP filter and the Miteq , 30 db gain) was added to the chassis box of the Angle linear ones.16 db pads added to output of it to equalize the gain with LRR 1.
ansanolla@naic.eduWednesday, March 2, 2005 16:32:20RG-9 coax cables #9 & 15 were replaced at the rotary joint, its jacket wre broken letting water to got unto them.Were replaced with RG-214 (the RG-9 replacement).Cable #9 is the 46.8 Mhz xmmiter link to CH #1 and #15 is the 430 Mhz xmitter pre-pulse. Also Heliax cables #31 & 33 were hooked to CH #1 for testing , they are getting there thru coax #4 & 10 respectely on RF cabinet 3 of CH#1.
ansanolla@naic.eduMonday, March 7, 2005 13:05:38Found that the the rf gating signal driver was not working cause the J. Mathews chassis ac power connector was loose and there is where the gating signal comes.Chassis in the need of a ac guard.
ansanolla@naic.eduMonday, March 14, 2005 09:44:22Dual beam run.Cooled amp been used.
ansanolla@naic.eduWednesday, March 30, 2005 15:11:33The 430 Mhz xmitter blew up the klystron that was replaced.So for a while the experiment run has been halted.
W. Genterwgenter@naic.eduTuesday, June 28, 2005 11:18:50On June 21 we arrived to find that there was a noise on the 430 receiver. This noise resembled the possible problem associated with the transmitter. Later when the transmitter was powered up the previous noise was always observed when the HV was above 30KV and was not present on lower HV. The pervious noise which we had noticed in the AM was not present. We investigated and found that in the CH we could hear what sounded like arcing coming from the Dewar Bias Cable. This noise was not present when the crosshead was off or when the bias supply is off. When the bias was on we could again hear the noise when tapping on the Dewar or input connectors with the compressor/crosshead off. Yesterday we installed the high level room temp amp and the transmitter was tested again (HV) this morning. Again when the HV was above 30 KV the offending noise was present once again. We reinstalled the Cryo Amps and tested again. The noise was present as before doing the same tests. The Bias supply was changed with the same result. Now the Cryo Amps are in the system and the original bias supply is reconnected. We suspect that if there is truly a problem, it resides in the Dewar. One more test to conduct and that is to completely remove the Dewar from the transmitter circuit and once again retest to see if the offending birdies reappear
ansanolla@naic.eduThursday, August 11, 2005 14:34:05System running a dual beam experiment. Is using the cooled amp.
ansanollla@naic.eduTuesday, August 16, 2005 11:33:18System after dual beam run was giving a wrong Tsys value at channel A. It was that the cal value was left for the radar people (181k) instead the RA people (47.5k)
ansanolla@naic.eduThursday, August 18, 2005 15:45:03Calixto , Carlos & Johanes were installing the John Matthew receiver box.
ansanolla@naic.eduFriday, August 26, 2005 13:09:11Due to power failure on 8/25/2005 system warmed up.
ansanolla@naic.eduTuesday, August 30, 2005 15:47:49System cooled . Back to normal operation.
ansanolla@naic.eduWednesday, August 31, 2005 15:42:57Calibrator of channel B was not switching properly.It had the 8T14 line driver bad on it.
ansanolla@naic.eduFriday, September 9, 2005 11:54:40Cal A attenuator driver was not working due to faulty 8T14 line receiver at the calibrator.
ansanolla@naic.eduTuesday, September 13, 2005 08:34:07Was running dual beam.Tsys normal after run 87 & 68k.
ansanolla@naic.eduWednesday, September 21, 2005 11:36:59Running dual beam . Tsys was checked and was 69 & 53 k.Also step att was working properly in both channels.
ansanolla@naic.eduTuesday, September 27, 2005 14:59:14System was recalibrated by WG. The new values are the following. Channel A 26.5k / 50.4k / 103.9k / 192.6k Channel B 15.8k / 33.0k /64.4k / 122.2k
ansanolla@naic.eduMonday, October 3, 2005 09:45:38System running dual beam. T sys is chann A= 88k & B=68k.
ansanolla@naic.eduThursday, October 6, 2005 08:07:21System back to normal on 10/4/2005.
ansanolla@naic.eduThursday, October 6, 2005 08:09:33System was used between 10/3/2005 to 10/4/2005 for dual beam.
ansanolla@naic.eduWednesday, October 12, 2005 10:42:26The Tsys data from this morning was showing that recvr was bad , pinpoint the problem to the downstairs LO drive chassis , it's power sw was off.
ansanolla@naic.eduMonday, November 7, 2005 14:14:15Dual beam run 11/11/2005.
ansanolla@naic.eduFriday, November 11, 2005 11:10:23Dual Beam run finished.
ansanolla@naic.eduThursday, November 17, 2005 11:50:00Dual beam enable.
ansanolla@naic.eduTuesday, December 27, 2005 11:09:13Dual beam run until 1/3/2006.
ansanolla@naic.eduThursday, January 12, 2006 14:56:21The 430 Mhz calibrator burned out its TTL line xmmiters & line receivers (8T13 & 8T14) from its cal value ckt.
ansanolla@naic.eduTuesday, January 31, 2006 13:48:48Yesterday we swap the 30 Mhz (10 Mhz B.W.) IF filter of the CRR2/LRR2 converter for FD for a 30 Mhz (20 Mhz B.W.)one.
ansanolla@naic.eduThursday, February 16, 2006 11:01:02The John Mathews 430 Mhz recvr was intermmitant , CR & BG found the ac line cord loose (corcom type).CR tied it up with cable straps to hold it in place.
ansanolla@naic.eduMonday, February 20, 2006 15:51:57John Mathews recvr still intermittant.Problem pinpointed to the rat race.
ansanolla@naic.eduTuesday, February 21, 2006 11:51:59The cryo comp of the 430 mhz Dewar replaced for maint. It was ED 022 (CTI), now is ED 022 (Austin).
ansanolla@naic.eduTuesday, February 21, 2006 11:54:22Mistake on last log , cryo comp now at CH #1 430 Mhz Dewar is ED 021 (Austin ).
ansanolla@naic.eduTuesday, February 21, 2006 11:55:44Rat race back into the John Mathews recvr.
ansanolla@naic.eduWednesday, April 5, 2006 11:58:07Cryo comp ED 021 (Austin) replaced by ED 022 (CTI).ED 021 is leaking.
ansanolla@naic.eduMonday, May 15, 2006 11:02:44John Mathew receiver removed.
ansanolla@naic.eduTuesday, May 23, 2006 13:26:38John Mathew reveiver back.Tsys is 300 k deg.
ansanolla@naic.eduWednesday, May 31, 2006 15:33:18Dual beam run since 5/27/2006 until 5/31/2006.
ansanolla@naic.eduMonday, June 19, 2006 08:54:11John Mathews receiver down for antenna relocation.
ansanolla@naic.eduMonday, June 26, 2006 11:21:04Dual beam run up until 7/5/2006.
ansanolla@naic.eduWednesday, July 5, 2006 15:45:36Dual beam ended.
ansanolla@naic.eduFriday, August 4, 2006 15:52:54The John Mathews recvr was installed to its new position. Also the new coax cables on it. It just improved the Tsys for 50 deg. Before we had a Tsys of 290 deg and now is around 250 deg.
ansanolla@naic.eduFriday, August 11, 2006 13:56:22BG measured the John Mathwes receiver upstairs.
ansanolla@naic.eduTuesday, August 22, 2006 14:42:24The LRR L.O. chassis (390-410 Mhz) was taken down to check an spike (420 Mhz) in the pass band.The 380m Mhz notch filter was retuned in order to lower down the 380 mhz spike in the pass band.Instead -30 db went to -46 dbm.
ansanolla@naic.eduThursday, August 24, 2006 16:22:30The CRR receivers L.O. was not working properly. Found a cable misplaced & an itermmitant one.
ansanolla@naic.eduTuesday, September 5, 2006 14:34:31Ganesh sent an e-mail tha the noise of the cryohead was not normal.We checked it upstairs a its noise is ths same that it was.Keep an eye on it.
ansanolla@naic.eduThursday, September 7, 2006 11:28:39John Mathews receiver gating sw limiter replaced with an sma type ACLM. It had the Alaskan one.
W. Genterwgetner@naic.eduThursday, September 14, 2006 09:44:03Two DCI cavity filters have been received for this receiver. They are centered at 430MHz with a BW of 20MHz.
ansanolla@naic.eduTuesday, October 17, 2006 15:46:59System Temp at channel A was too high.It was pinpointed to the receiver calibrator at CH 1.It had both line receivers burned out (8T14). Weinchel att. were not switching
ansanolla@naic.eduWednesday, October 18, 2006 16:13:09New L.O./R.F. gating sw / 400-460 Mhz sw chassis installed at CH 1 , also it will drive the Dome gating signal thru f.o. cable.
ansanolla@naic.eduFriday, October 20, 2006 11:44:34Dual beam on.New gating chassis on duty.
ansanolla@naic.eduWednesday, November 1, 2006 09:19:46The 430 Mhz monoplexer was taken down for measurments . So a 50 ohm termination is in front of the receiver dewar at channel A. Operations was notified.
ansanolla@naic.eduThursday, November 2, 2006 11:39:28Channel A back to service after replacing the monoplexer.T sys is 84 k deg , B is 58 k deg.Monoplexer driver still with problems.
ansanolla@naic.eduFriday, November 17, 2006 11:27:50System in dual beam operation.
ansanolla@naic.eduTuesday, December 5, 2006 15:43:24Forgot to mention that dual beam operation was removed since 11/22/2006.Now on RA mode.
ansanolla@naic.eduFriday, April 20, 2007 11:13:28Cryo remote reading was not working properly due that the Lakeshore meter was left on. Mea culpa.
ansanolla@naic.eduMonday, April 23, 2007 09:45:02System down until further notice due to painting project.
ansanolla@naic.eduWednesday, April 25, 2007 13:04:23System cryo comp taken down for service.
CRVcrios@naic.eduWednesday, January 2, 2008 14:22:49Vacuum pump conected at dewar and is pumping.
crvWednesday, January 30, 2008 14:32:19System started cooldown on 1/03/2008. The system is cold.
crvWednesday, April 30, 2008 14:03:07Limiter installed to post amp.Chanel A output.
crvWednesday, May 7, 2008 09:25:59Limitter change to the postamp. CHA output was actualized in the auto cad drawing.
crvThursday, July 31, 2008 16:36:08problem with the calibration. Cal chassis removed and checked Ics in driver board where re-inserted in ic sockets. Chasis was installed and system is working fine.
DEWdwhitlow@naic.eduTuesday, April 21, 2009 07:41:49Replaced CHA LNA in dewar with experimental uncooled amplifier, to see if radar recovery problem is solved or at least reduced significantly. Change was made approx 1500 AST, 20 April 2009.
crvThursday, August 6, 2009 15:07:26reciever was vacuumed and is cooled after sunday night power failure and wednesday power failure .
crvThursday, August 20, 2009 15:02:38bais box removed from dewar on tuesday august 18 2009 by Dana and installed a bias box desingned by him. Temp can't be measured and tsys will have to be used as reference to know if reciever is warming up.
Dana Whitlowdwhitlow@naic.eduFriday, August 21, 2009 15:49:1418 Aug 2009: replaced bias chassis with experimental battery-powered open-loop bias box to eliminate ground loops and hopefully improve RF overload recovery behavior. Initial battery voltage (no load) = 6.5V. Initial battery voltage under load = 6.23V. 21 Aug 2009: battery voltage under load 5.59V. Feeling is that it's running down somewhat faster than expected (isn't that always the case?)
Dana Whitlowdwhitlow@naic.eduThursday, September 3, 2009 14:21:49Updates on battery voltages: 27 Aug: 5.19 V (Vd battery) 3 Sept: 4.5 V (Vd battery) It's time to swap the batteries or get some power supplies!
crvTuesday, September 8, 2009 14:05:09Dana's bais box removed and original Bias box installed. Cryo temp. can now be read from control room
Dana Whitlowdwhitlow@naic.eduTuesday, September 8, 2009 14:29:21Radar test conducted 8Sept09 with original bias system shows that the recovery problem has returned, adding confirmation to the impression that the floating, open-loop bias box is the key to much better recovery behavior.
crvcrios@naic.eduMonday, March 22, 2010 16:05:48Channel A dewar input terminated with 50 ohms and momoplexer output terminated with 50 ohms
GRgrivera@naic.eduWednesday, June 30, 2010 16:13:10On Tuesday, June 29,2010, the crosshead ED015 was replaced with ED038 (straight from the XB Rx, that was in the Lab for testing). Today, June 30, the cool down started.
crvTuesday, September 28, 2010 11:35:49Receiver amp. were tested and found that Ch A amp. has a short on stage 2. The cryopump has been turned off and will be serviced tomorrow after it warms up completely.
crvcriosnaic.edu@Thursday, September 30, 2010 16:34:39The amp in Ch A with serial number 450-13 was replaced by 450-18 system started cool down.
crvFriday, October 1, 2010 11:42:15System is cold.
crvFriday, October 1, 2010 11:44:38problem in the signal path was detected on september 27th 2010 has been solved . There was a cable misconnected.
crvcrios@naic.eduMonday, December 13, 2010 11:59:51the dewar was warmed up on friday because we found it condensating and the cryo temp was 179K. This was a effect of a power failure. The system was vac. and started cooldown on saturday Dic.11th by Dana today we closed the valved and connected Dana's Bias box.
criosMonday, June 6, 2011 14:05:51tsys on ch A were bad neg values. the problem was on chasis R166 on the control room. IC 9,10,11,13 were replaced (mc3487) tsys are ok.
Dana Whitlowdwhitlow@naic.eduWednesday, November 23, 2011 17:01:24The 430CH RX was recently brought down to replace a broken fridge. While it was down we investigated a long-standing problem of apparent poor Tsys on CHA, which as been running about 35 to 40K higher than that of CHB. A check with the NFA showed that both LNAs were well-matched to each other, with Tn values around 24K. But following reinstallation in the dome, the CHA Tsys was found to be running almost 75K higher than that of CHB- the problem was now much worse! This was investigated this morning, and it seems like the Jennings switch is at fault. After playing with it, the Tsys difference has fallen to a much more reasonable value of ~15K. Some difference is to be expected because of the presence of the monoplexer in the CHA path. We plan to outright replace the Jennings switch Monday morning (5 days hence).
Dana Whitlowdwhitlow@naic.eduWednesday, June 27, 2012 16:53:47Measured Gain and Tn on newly refurbished CH receiver: Dewar temp: ~17K (by VPG) CHA: Vd = 1.50V Id1 = 4.8 mA Vg1 = -1.790 V Id2 = 4.5 mA Vg2 = -1.776 V Tn = 18.8K Gain = 23.5 dB CHB: Vd = 1.50V Id1 = 4.8 mA Vg1 = -1.781 V Id2 = 4.8 mA Vg2 = -1.766 V Tn = 18.3K Gain = 24.5 dB Measurements were made thru swept SMA elbow and SMA-N adapter (neither included in NFA calibration). Measurements were made using the old 346A noise head (could not find the new head).
Denis Urbaindurbain@naic.eduTuesday, September 11, 2012 13:02:17The Tsys was giving wrong reading, the 260MHZ output was not seen on the spectrum analyzer. The 30-260 converter was not getting the right LO frequency from the synthesizer. Setting the synthesizet to local and the Frequency by hand to 230MHz or 290MHz gave correct Tsys. Problem was resolved by Phil doing a reset of the Synthesizer.
Denis Urbaindurbain@naic.eduMonday, December 17, 2012 16:15:23After the dual beam run, the Pol A Tsys was much higher than the Pol B. Before the dual beam both channels had similar temperatures. Initial Tsys gave Pol A:73K and Pol B:58K The DPDT RF switch was operated several times and each time the Tsys was measured. The results were the following: 140/58, 112/55, 81/61. It seems that the contacts inside the switch are not repeatable. Microwave Techniques the original manufactuer doesn't exist anymore, the company was acquired by MegaRF solutions. Conacted MegaRF to purchase new copper contacts.
Denis Urbaindurbain@naic.eduFriday, May 24, 2013 11:05:49Power outage on the 18 of May left the compressor off and the receiver started to warm-up. Since the receiver was not needed immediately, we left it warm up completely and Guarionex serviced the cold head (baking of the displacer and reinstalling the 350 motor instead of the 1020). The receiver started cooling on the 23rd.
Dana Whitlowdwhitlow@naic.eduWednesday, November 13, 2013 17:17:11Today we went looking again for the problems with pol A of the line feed receiver: generally excessive Tsys, and recently-reported severe gain changes. We radiated a CW signal from the large log-periodic calibration "dipole" antenna, and viewed it in the CH with the R&S ZVL SA, through the entire upstairs signal path. The intent was to bang on things looking for a response. The intent was somewhat corrupted by the fact that we could not rotate the AZ for the purpose of pointing the line feed close to the CAL dipole. When we tried rotating the CAL dipole, I noticed that the system seemed to exhibit a large axial ratio of ~10 dB. We ran the CH down to 0 ZA, and the axial ratio improved to ~2 dB, but not monotonically at all. And the signal strength only improved by ~3 dB. So it looks likely that the footprint of the line feed may have large areas of really bad circularity- unless, of course, the CAL dipole has big problems. But I digress. Back to the original problem, I also noticed that the received signal had AC line-locked negative going spikes in amplitude, which was eventually (and rather painfully) traced back to the 20-MHz source downstairs in rack 3. This was causing severe dropouts in the L.O. signals to the downconverters in the CH. I have requested a technician to take a look at the downstairs source- I suspect a bad power supply. We never did get to banging on things in the CH, reasoning that the power supply problem could easily obscure transient events (and we were getting tired of the CH, after too much time spent tracing cables in all that mess).
Dana Whitlowdwhitlow@naic.eduThursday, November 21, 2013 16:10:09The 20MHz source (Chassis 3-4, downstairs) has been repaired. As thought, the problem was a failing power supply, in this case the 15V supply. The IF passband is free of 120Hz dropouts, and the amplitude margin on both 20MHz outputs is at least 7 dB.
Dana Whitlowdwhitlow@naic.eduFriday, November 22, 2013 11:51:47Felix and crew went to the CH and mechanically "stimulated" the various elements in the signal path in pol A, whilst watching the signal level of a CW tone injected from the CAL dipole. No change whatsoever was reported. While the above was going on, I had a chance conversation with Phil, who suggested that perhaps the chronically-high Tsys values could be a CAL problem. I rechecked Tsys, and found only a 15K differential between the pols (probably a proper healthy value) instead of the ~40K noted earlier this morning before they performed the aforementioned stimulations. Felix told me that in the course of the stimulations, they did a round of connector tightness checks, which could have affected either the signal path or the CAL path. Anyway, now we watch and wait some more!
Dana Whitlowdwhitlow@naic.eduFriday, May 9, 2014 10:06:46The problem of Pol A's Tsys being grossly higher than that of Pol B has returned, apparently as of about May 7 2014. The CAL switches in the radar equipment room have been checked and are set correctly. I'm told that seemingly unrelated work (turning a crosshead back on) was done in the CH at vaguely about the right time- perhaps a cable was snagged (or something) during the process.
CRR1 not visible downstairsfelix.fernandez@naic.eduMonday, September 22, 2014 07:28:50On 18-SEP-2104 during T2879 run CRR1 was found to be not responding to load or ant by Nestor Aponte. After some debugging of R107 inside the carriage house we concluded that none of the two amplifiers within CRR1's path was at fault the only likely candidate was thought to be the mixer. We tried replacing CRR1's mixer (NOTE: to accommodate the new mixer we had to switch the lids of the mixer. Hence, the labeling of LO and RF are now swapped.) and had some minimal improvement upstairs but the problem persisted downstairs. After tracing the cable Heliax #1 we found out that the 6 port coax switch was stuck in some other port, not port #2 as it should have. We connected Heliax #1 straight onto CRR1 output using and female-female N-type connector. This should be a temporary fix until we can replace the 6 port coax switch.
Bad LRR LOfelix.fernandez@naic.eduWednesday, October 8, 2014 15:00:56The LRR paths were missing their LO signals. The problem was due to a faulty BNC connection on Coax #13 (Gated 20MHz signal) which was attenuating the signal by 10dB. The attenuated 20MHz signal is a required input for R-598 which generates the LRR's LO signals.
Bad Tsys valuesfelix.fernandez@naic.eduWednesday, October 8, 2014 15:03:36After repairing the LRR's LO the Tsys values were suspicious. The nominal values are >80K and the Tsys we were getting were <30K. The problem was the line drivers (8T13) in Rack 8-1 inside the control room. The line drivers were dead and once replaced all Tsys were nominal.
Garbage Tsys readings (Rack 106)felix.fernandez@naic.eduFriday, December 12, 2014 16:10:47Today's morning Tsys readings of PolA 430CH were atypical. In fact, the readings made no sense. After debugging the system the problem was found to be in R106. R106 is 430CH's diode/attenuator driver box. We brought R106 for further inspection. The only thing we noticed was a loud clicking noise when we R106 was powered-up. The sound seemed to come from the step-attenuator. Once R106 was replaced this afternoon the Tsys readings were back to normal.
Cold-head maintenancefelix.fernandez@naic.eduTuesday, December 30, 2014 15:49:13Due to ratcheting sounds the 430CH's cold-head was serviced today. The cold-head was removed, cleaned, but not "baked". This procedure should buy us some extra time for next weeks' SB and DB. Complete servicing (i.e. baking the cold-head's displacer) should be done soon.
Faulty Comp A Coax Swithcfelix.fernandez@naic.eduThursday, January 15, 2015 11:55:29Yesterday's afternoon Tsys readings had erroneous values on Ch-A of 430CH. After debugging we found out that Comp A coax switch was not resetting to the desired position. After replacing the coax switch all Tsys values are back to normal.
Noise in the cryopumpcrios@naic.eduThursday, July 9, 2015 12:57:22 The cryopump had a noise that was similar to contamination. We changed the power supply and the noise stop. System is cooling fine.
Cryo pump changecrios@naic.eduTuesday, September 22, 2015 13:53:57Cryo pump turned off on sunday september 20th do to contamination.Cryopump replaced on monday 21st and started cooling . Vac. pump removed today tuesday 22nd system ready to use.
Pol-B ANT/LOAD switch damagedfelix.fernandez@naic.eduTuesday, November 17, 2015 11:25:15Pol-B Tsys reading were stuck at 300K, implying that the system was looking at LOAD instead of SKY. After some debugging the relay switch between ANT/LOAD on Pol-B was found to be damaged. It will be replaced as soon as possible.
Cold-head knocking soundfelix.fernandez@naic.eduTuesday, February 2, 2016 14:54:13The 430CH's cold-head started "knocking". We started to warm-up the receiver today (Feb-2-2016) and will replaced the cold-head tomorrow.
Compressor trip/resetfelix.fernandez@naic.eduFriday, June 2, 2017 13:15:02Since 5/30 the temperatures on the 430CH receiver were not logged. Today, upon noticing that the CH receiver temp was reporting 300K the CH was inspected and the compressor was in the OFF state. After resetting the compressor breaker (which is the same ON/OFF switch) the compressor was turned back on. Now we only wait for the Dewar to evacuate so that we may start the cool-down process.


Back to Top